ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: 17, 81, 82 Fring, Andreas, Freie Universität Berlin, Berlin, Germany Lie algebraic Structures from factorized Virasoro Characters Integrable deformations of certain conformal field theories are known to exhibit a Lie algebraic structure related to affine Toda field theories. Exploiting the fact that the related characters factorize, we demonstrate that precisely this structure is already present in the conformal field theory. We employ these results in the construction of the underlying statistics and quasi-particle states. We also discuss some implications on the thermodymamic Bethe Ansatz. Ref. A. G. Bytsko and A. Fring, Thermodynamic Bethe Ansatz with Haldane Statistics. A. G. Bytsko and A. Fring, Lie algebraic Structures from factorized Virasoro Characters.